if the O2 is reacting at the rate of 0.25 Ms-1,


calculate:


i. rate of NO2


1
Expert's answer
2022-03-19T02:44:56-0400

Consider,

2NO(g)+O2(g)→2NO2(g)

Moreover, understand that reaction rates are proportional to the stoichiometry of the reaction. Hence,


"\frac{0.070M}{S} \times \frac{2NO_2}{2NO}=\frac{0.070M}{S}"


is the rate of appearance of nitrogen dioxide, and


"(0.070\frac{M}{s})\times \frac{0_2}{2NO}=\frac{-0.035 M}{s}"


is the rate of disappearance of the oxygen.

To be sure, reaction rates are merely,

"\frac{\bigtriangleup[A]}{\bigtriangleup t}"

and the concentrations of the constituents are proportional to the stoichiometry.